# Exploring run-and-tumble movement in confined settings through simulation

### v1

@@ Abstract
Motion in bounded domains is a fundamental concept in various fields, including billiard dynamics and random walks on finite lattices, with important applications in physics, ecology and biology. An important universal property related to the average return time to the boundary, the Mean Path Length Theorem (MPLT), has been proposed theoretically and confirmed experimentally in various contexts. In this discussion, we investigate a wide range of mechanisms that lead to deviations from this universal behavior, such as boundary effects, reorientation and memory processes. In particular, this study investigates the dynamics of run-and-tumble particles within a confined two-dimensional circular domain. Through a combination of theoretical approaches and numerical simulations, we validate the MPLT under uniform and isotropic particle inflow conditions. The research demonstrates that although the MPLT is generally applicable for different step length distributions, deviations occur for non-uniform angular distributions, non-elastic boundary conditions or memory processes. These results underline the crucial influence of boundary interactions and angular dynamics on the behaviour of particles in confined spaces. Our results provide new insights into the geometry and dynamics of motion in confined spaces and contribute to a better understanding of a broad spectrum of phenomena ranging from the motion of bacteria to neutron transport. This type of analysis is crucial in situations where inhomogeneity occurs, such as multiple real-world scenarios within a limited domain. This bridges the gap between theoretical models and practical applications in biological and physical systems since the study of the statistics of movement in confined settings can bring some light in explaining the mobility mechanism of active agents.
1 Introduction
The properties of motion within bounded domains are applicable across a variety of fields such as physics, ecology and biology. Motion in bounded domains has far-reaching implications, from fundamental mathematical insights to practical applications in environmental science, biology, medicine and ecology. It lays the foundation for predicting and controlling the behavior of particles or agents in diverse and restricted environments, offering potential benefits to numerous scientific and technological fields.
A key aspect of motion in confined environments is the Mean Path Length Theorem (MPLT), a generalization of a geometric result originally proposed by Cauchy [1]. The (MPLT) is a fundamental result in the study of particle dynamics in confined domains, providing a universal expression for the average path length of particles. In particular, this theorem states that the mean path length $<L>$ of particles within a bounded domain $\Omega$ is determined by the ratio between the volume $V_{\Omega}$ and the surface area $\Sigma_{\Omega}$, regardless of the details of the dynamics of the particles or the shape of the domain.
@@ 1 Introduction
where $\eta_{d}$ is a numerical factor depending only on space dimensionality $d$ (in the case we will consider in detail $d=2$ and $\eta_{2}=\pi$).
@@
More precisely, the MPLT concerns the case where $<L>$ is the mean length of randomly distributed chords intersecting a convex body. Based on the linear Boltzmann equation, the MPLT was also extended to scenarios in which the particles do not move in straight lines but are scattered [2]. This showed that the Cauchy formula applies more generally to the stochastic paths of Pearson random walks with exponentially distributed flight lengths. This demonstrates that under the conditions of a uniform and isotropic particle influx, the absence of absorption or particle reproduction within the medium, and a detailed balance in scattering probabilities, the average path length remains unchanged. The analysis contributes to the understanding of transport phenomena in scattering media, which are relevant, for example, for the estimation of reaction rates in neutron transport [3, 4, 5], ionic recombination in gasses and other processes in finite volumes. Furthermore, this theorem has recently been generalized for non-convex domains and extended to the case of Brownian motion [6, 7, 8]. The work of Benichou et al. [9] discusses the averaged residence times of stochastic motions within bounded domains, focusing on the mean first exit and residence times for particles undergoing random motions, such as Pearson random walks, by employing a backwards Chapman Kolmogorov equation. This study is motivated by various physical systems, including animal trajectories [10] and neutron scattering processes. In [11] the case of arbitrary closed trajectories in arbitrary domains is investigated and it is shown that the mean path length still agrees with the Cauchy formula as long as no closed trajectory is completely contained in the domain. Shukla et al. [12] extend the Cauchy formula to lattice-based random walks, providing a straightforward understanding compared to previous models relying on continuum-based Boltzmann transport equations. By analyzing random walks on an $N\times N$ square lattice, the research provides both numerical results and theoretical underpinning of the observed phenomena, suggesting that the Cauchy formula can be generalized beyond traditional applications to include more complex random walk scenarios. This property holds for anomalous and classical photon propagation [13]. All these generalizations imply that the Cauchy formula, Eq. (1) is applicable to a wide range of scenarios and phenomena, hence the importance. Cauchy’s universality was recently confirmed by experimental results on the movement of bacteria [14] and the propagation of light in scattering media [15, 16].
@@
In addition, to know the statistics of random walks in confined settings offer insights into a range of phenomena, such as the conditions under which active particles can escape barriers or be trapped by external forces, or how chemotactic fields influence particle motion towards or away from high nutrient concentrations, applicable to bacterial motion, intracellular transport, and animal foraging, bridging the gap between theoretical models and real-world biological processes [17, 18, 19]. These systems presents what is called a Run-and-tumble movement, characterized by randomly reoriented ballistic movements. Understanding the mean chord length can provide foundational insights into the geometry and scale of motion within a domain, which is indirectly relevant to estimating escape times or paths in narrow escape scenarios, especially when considering the impact of domain geometry on escape dynamics [20, 21, 22] or target searching [23].
@@
It is widely recognized that numerous stochastic transport phenomena observed in the real world do not exhibit jump lengths that are exponentially distributed, nor do they show scattering angles that are uniformly random [25, 26, 27]. For instance, isotropic random flights having fixed length have application in polymer chains [28] and an exponential law has application, as mentioned, to neutron diffusion processes and locomotion problems in biology. It naturally raises the question of whether these processes still adhere to similar universal characteristics. In fact, heuristic reasoning indicates that Cauchy’s formula is applicable to any (isotropic) random walks as long as the random walk’s entry into the domain involves a length distribution that aligns with equilibrium [9, 29, 30]. In particular, some attention was recently given in the literature to the role of boundary conditions in influencing the absorption time of run-and-tumble particles [31, 32], and to the particle’s speed distribution or the spatial dimensionality [33].
@@
Complex dynamics in bounded domains are occasionally linked to non-homogeneous distributions, as in the case of active particles (see, for example, [34, 35, 36]) or when interactions with walls are present [37]. Therefore, it is natural to wonder if Cauchy universality is upheld in these situations. This is the main topic we will cover, giving examples of mechanisms causing nonuniform stationary probability distributions and breaking Cauchy universality; we will take into account both the scenario where particles move in a stochastic way inside the body (random walk) and the scenario where they follow rectilinear trajectories between consecutive collisions with the boundary (billiard). Through theoretical analysis and numerical simulations, this study examines how different boundary and reorientation conditions, affect the motion’s universal properties. This research offers potential implications for understanding motion in confined systems across various scientific domains, extending beyond the classical applications of billiards and random walks to include biological and physical phenomena where such conditions are prevalent.
@@
The paper is structured as follows. In the next section, the methodology for simulating particle movement within a 2D circular space is described, including variations in step length and angular distributions under different boundary conditions. Section 3 presents the analysis confirming the MPLT under uniformly random angular distributions and highlights deviations with non-uniform distributions. The section 4 examines path length distributions and their variance, illustrating the transition from diffusive to quasi-ballistic movements. Section 5 explores the radial probability distributions, showing uniformity under elastic boundary conditions and deviations with random angle conditions. In section 6, we introduce memory effects in particle reorientation and their impact on path length and radial distributions. Finally, we summarize the key findings and their implications in section 7.
2 Simulations
@@
Figure 1: Trajectory of the particle inside a circle with elastic boundary. Different types of movement can be obtained by changing the ratio between the mean step length and the circle radius, $l/R$: (a) $l/R=0.01$, (b) $l/R=0.1$, (c) $l/R=1$, (d) $l/R=10$. Figure (a) is diffusive, on the other extreme, (d) is quasi-ballistic.
@@
To replicate a simple geometric setting common to experimental realizations [18, 19, 14, 37], our code simulates a particle moving within a unit 2D circular space. In our model, the particle travels at a constant unit speed, $v=1$. Therefore, the time between successive collisions is equal to the total trajectory length, allowing us to discuss average residence time and mean path length interchangeably. In two dimensions, the mean path length, as defined in eq. (1), is expressed as follows:
@@
where $l=<\ell>$ is the mean step length and $\sigma$ is the standard deviation, (b) uniformly random, (c) fixed (as in polymer chains [28]), and (d) exponential distributed (typical of bacteria mobility [14] and neutron scattering [3]),
@@
In the case of the Gaussian distribution, no cutoff was used. This means that there is a possibility that the particles move backwards if $\ell$ takes a negative value. We will analyze the statistics and the validation of the MPLT under various angular distributions: (a) a fixed angle, which can be considered as a representation of strong chemotaxis (preferred direction); (b) a Gaussian distribution, representing weak chemotaxis, since there is a preferred direction with some deviation, a similar approach is used in [38]; (c) the particle is allowed to move in any direction uniformly distributed in $[0,2\pi]$, which is the typical case considered in most studies.
@@
By adjusting the mean step length, $l$, we can observe different types of movement, as illustrated in Fig. (1). While these examples start from the center of the circle, the findings presented in this work are not dependent on the starting location, including points on the boundary. In Fig. (1), the exit angle after each tumble event is uniformly random, and the step length follows an exponential distribution. We observe various movement types according to the ratio $l/R$, where $R$ is the circle’s radius. When $l/R<<1$, the movement is diffusive, but when $l/R>>1$, the movement becomes quasi-ballistic. Purely ballistic movement occurs in the limit as $l/R\to\infty$.
3 Mean path length theorem
First, we explore the conditions under which the Mean Path Length Theorem (MPLT) is valid. To do this, we calculate the mean path length using various combinations of step length and angular reorientation distribution. This analysis is conducted with elastic boundary conditions for four distinct values of $l/R$: $0.01$, $0.1$, $1$, and $10$. We find that the MPLT is valid only when the angular distribution is uniformly random, across all step length distributions. At first, the fact that the mean path length is the same for different values of $l/R$ seems very surprising, because if we replace straight segments (or arcs) by irregular paths, we intuitively expect a twofold mechanism to profoundly change the dynamics (stochastic short returns to the boundary and long walks in the domain without touching the boundary) [29]. This fact is confirmed by plotting the ratio $(<L>_{s}/<L>_{t})-1$, where $<L>_{s}$ represents the mean path length from the simulation, and $<L>_{t}$ is the theoretical value predicted by the MPLT. If the MPLT is accurate, the simulated and theoretical mean path lengths will converge as the number of collisions increases. Here, we present some of the most illustrative cases. In Fig. (2), a typical graph of $(<L>_{s}/<L>_{t})-1$ is shown for various $l/R$ ratios, using an exponential step length distribution and a uniformly random angular distribution. It is observed that the difference between simulated and theoretical values tends to zero as the number of collisions increases.
@@ 3 Mean path length theorem
Figure 2: Difference between predicted value of the mean path length according to the theorem and the value of the simulation versus the number of collisions with elastic boundary condition, exponential step length distribution, and uniformly random angular distribution. Each point is an average of 20 simulations. It can be seen that they coincide when the number of collisions approach to infinity.
@@
The step length distribution can be adjusted to any of the mentioned scenarios, and the MPLT will still hold true. For illustration, in Figs. (3(a)) and (3(b)), we present two distinct cases of step length distribution: (a) random between 00 and $2l$, and (b) a constant value of $l$. The differences between the simulated and theoretical values of mean path length approach zero as the number of collisions tends towards infinity in all instances.
@@
Figure 3: Difference between predicted value of the mean path length according to the theorem and the value of the simulation versus the number of collisions with elastic conditions at the boundary, random angle distribution, and (a) random step length between 0 and $2l$, (b) constant step length $l$.
On the other hand, if we alter the distribution of the reorientation angle after the tumble events, the assumption of homogeneous flux at the boundary no longer applies, and consequently, the mean path length does not converge to the same value for all $l/R$ ratios, as illustrated in Figs. (4(a)) and (4(b)). In Fig. (4(a)), the angles after the tumble event follow a Gaussian distribution with a mean of 00 and a standard deviation of $\pi/2$. We observed similar outcomes with various combinations of means and standard deviations. Similarly, Fig. (4(b)) displays results for a fixed angle of $\pi/4$, yet comparable results are achieved with different preferred direction values.
@@
Figure 4: Difference between predicted value of the mean path length according to the theorem and the value of the simulation versus the number of collisions with elastic conditions at the boundary, exponential step length distribution, and (a) gaussian angle distribution with mean 0, and $\sigma=\pi/2$; (b) fixed value of angle at $\pi/4$.
The fixed angle and the Gaussian distribution around an angle create a preferred direction, observable in some experimental behaviors of bacteria due to mechanisms like chemotaxis. The fixed angle constitutes a strong condition (the particle consistently moves in a specific direction), while the Gaussian distribution signifies a less stringent condition (there is a preferred direction, yet a certain probability exists for the particle to move in another direction close to the preferred one). In scenarios with fixed and Gaussian distributed angles, the particle tends to collide with the boundary consistently in the same region, as illustrated in Fig. (5). This results in non-uniformity and anisotropy, thus violating the assumptions underlying the MPLT. This result is not surprising, as previous works had already clarified the conditions of the angular distribution under which the theorem holds [2, 7, 9, 30].
@@
Figure 5: Trajectory of 1000 steps of the particle inside a circle with elastic boundary, exponential step length distribution, and $l/R=0.1$. (a) The angle distribution is normal distributed around $\theta=0$, (b) Angle distribution fixed at $\pi/4$. The particle has a strong tendency to moves always to the same part of the border, producing anisotropy and therefore breaking the hypothesis of the MPLT.
Let’s now examine the scenario where the outgoing angle after a collision with the boundary is random. Initially, the absolute values of the mean path length under elastic boundary conditions, with an exponential step length distribution and uniformly random angle distribution, are presented in Fig. (6(a)). These values approach the theoretical value of $\frac{\pi}{2}=1.57079...$ as the number of collisions approaches infinity. For contrast, the absolute values of the mean path length for the random reflecting angle condition are plotted in Fig. (6(b)). As can be seen, the invariant property does not hold anymore, as we anticipated in [32]. On the contrary, the mean path length does depend on $l/R$, i.e., the type of movement (diffusive or ballistic), as can been appreciated in Fig. (6(b)).
Figure 6: (a) Mean path length versus the number of collisions with elastic reflecting conditions at the boundary. Each point is an average of 20 simulations. It can be seen that they approach to the theoretical value as the number of collisions tends to infinity. (b) Mean path length versus the number of collisions with random reflecting conditions at the boundary. Each point is an average of 20 simulations. In this case, the mean path length in not invariant and depends on the value of $l/R$.
In Fig. (7), we show this dependency in a $<L>$ versus $l/R$ plot for simulations with $10^{9}$ collisions. It can be seen that the mean path length tends to $4/\pi$, which is the mean path length of a random billiard as showed in [32].
Figure 7: Mean path length as a function of the parameter $l/R$ for simulations with $10^{9}$ collisions and random angle boundary conditions. The limit for $l/R\rightarrow\infty$ is the mean path length of a random billiard, $1.27324/\pi=$ 1.2732 ….
The MPLT holds true across various step length distributions -whether Gaussian, uniformly random, fixed, or exponentially distributed- as these do not affect the uniformity of particle incidence on the boundary, a critical assumption of the theorem. However, altering the angular distribution of the tumble events or the angle post-collision breaks this assumption. When the angular distribution deviates from being uniformly random, it introduces anisotropy or non-uniformity in the way particles interact with the boundary, violating the MPLT’s conditions. For instance, fixed or Gaussian-distributed angles after tumble events create preferred directions of movement, leading to non-homogeneous wall collisions and thereby the theorem’s failure. The MPLT does not hold under random boundary conditions, as opposed to elastic conditions, due to the way these conditions alter the uniformity and isotropy of particle interactions with the boundary of the domain. On the contrary, in elastic boundary conditions, a particle’s angle of reflection equals its angle of incidence, preserving the distribution of angles at which particles strike and leave the boundary. This uniformity ensures that the particle flux across the boundary maintains the assumptions required for the MPLT to apply.
4 Distributions of path length between collisions
It is evident that under the condition of homogeneous and isotropic flux of particles at the boundary, the mean path length is determined by Cauchy’s formula. However, it prompts the question: how does the distribution behaves in different conditions? Let us delve into this inquiry.

### v2

@@ Abstract
Motion in bounded domains is a fundamental concept in various fields, including billiard dynamics and random walks on finite lattices, and has important applications in physics, ecology, and biology. An important universal property related to the average return time to the boundary, the Mean Path Length Theorem (MPLT), has been proposed theoretically and experimentally confirmed in various contexts. We investigated a wide range of mechanisms that lead to deviations from this universal behavior, such as boundary effects, reorientation, and memory processes. This study investigates the dynamics of run-and-tumble particles within a confined two-dimensional circular domain. Through a combination of theoretical approaches and numerical simulations, we validate the MPLT under uniform and isotropic particle inflow conditions. The research demonstrates that although the MPLT is generally applicable for different step length distributions, deviations occur for non-uniform angular distributions, non-elastic boundary conditions, or memory processes. These results underline the crucial influence of boundary interactions and angular dynamics on the behavior of particles in confined spaces. Our results provide new insights into the geometry and dynamics of motion in confined spaces and contribute to a better understanding of a broad spectrum of phenomena ranging from the motion of bacteria to neutron transport. This type of analysis is crucial in situations where inhomogeneity occurs, such as multiple real-world scenarios within a limited domain.
preprint: AIP/123-QED
I Introduction
The properties of motion within bounded domains are applicable across a variety of fields such as physics, ecology, and biology. Motion in bounded domains has far-reaching implications, from fundamental mathematical insight to practical applications in environmental science, biology, medicine, and ecology. This lays the foundation for predicting and controlling the behavior of particles or agents in diverse and restricted environments, offering potential benefits to numerous scientific and technological fields.
A key aspect of motion in confined environments is the Mean Path Length Theorem (MPLT), a generalization of a geometric result originally proposed by Cauchy Cauchy (1850). The (MPLT) is a fundamental result in the study of particle dynamics in confined domains, providing a universal expression for the average path length of particles. In particular, this theorem states that the mean path length $<L>$ of particles within a bounded domain $\Omega$ is determined by the ratio between the volume $V_{\Omega}$ and the surface area $\Sigma_{\Omega}$, regardless of the details of the dynamics of the particles or the shape of the domain.
@@ I Introduction
where $\eta_{d}$ is a numerical factor that depends only on the space dimensionality $d$ (in this case, we consider in detail $d=2$ and $\eta_{2}=\pi$).
@@
More precisely, the MPLT concerns the case in which $<L>$ is the mean length of randomly distributed chords intersecting a convex body. Based on the linear Boltzmann equation, the MPLT was also extended to scenarios in which the particles do not move in straight lines but are scattered Bardsley and Dubi (1981). This indicates that the Cauchy formula applies more generally to the stochastic paths of Pearson random walks with exponentially distributed flight lengths. This demonstrates that under the conditions of a uniform and isotropic particle influx, the absence of absorption or particle reproduction within the medium, and a detailed balance in scattering probabilities, the average path length remains unchanged. The analysis contributes to the understanding of transport phenomena in scattering media, which are relevant, for example, for the estimation of reaction rates in neutron transport de Kruijf and Kloosterman (2003); Dirac (1943); Case (1953), ionic recombination in gases and other processes in finite volumes. Furthermore, this theorem has recently been generalized for non-convex domains and extended to the case of Brownian motion Blanco and Fournier (2003); Mazzolo (2004); Santalo (1976). The work of Benichou et al. Benichou et al. (2005) discusses the averaged residence times of stochastic motions within bounded domains, focusing on the mean first exit and residence times for particles undergoing random motions, such as Pearson random walks, by employing a backwards Chapman Kolmogorov equation. This study is motivated by various physical systems, including animal trajectories Jeanson et al. (2003) and neutron scattering processes. The case of arbitrary closed trajectories in arbitrary domains was investigated by Caballero et al. Hidalgo-Caballero et al. (2022), and it is shown that the mean path length still agrees with the Cauchy formula as long as no closed trajectory is completely contained in the domain. Shukla et al. Shukla and Thongjaomayum (2019) extend the Cauchy formula to lattice-based random walks, providing a straightforward understanding compared to previous models relying on continuum-based Boltzmann transport equations. By analyzing random walks on an $N\times N$ square lattice, the research provides both numerical results and theoretical underpinnings of the observed phenomena, suggesting that the Cauchy formula can be generalized beyond traditional applications to include more complex random walk scenarios. This property holds for anomalous and classical photon propagation Binzoni and Martelli (2022). All these generalizations imply that the Cauchy formula, Eq. (1), is applicable to a wide range of scenarios and phenomena, hence the importance. Cauchy’s universality was recently confirmed by experimental results on the movement of bacteria Frangipane et al. (2019) and the propagation of light in scattering media Pierrat et al. (2014); Savo et al. (2017).
@@
In addition, knowing the statistics of random walks in confined settings offers insights into a range of phenomena, such as the conditions under which active particles can escape barriers or be trapped by external forces, or how chemotactic fields influence particle motion towards or away from high nutrient concentrations, applicable to bacterial motion, intracellular transport, and animal foraging, bridging the gap between theoretical models and real-world biological processes Angelani, Leonardo, and Paoluzzi (2014); Rashid, Long, and Singh (2019); Long et al. (2017). These systems present what is called a Run-and-tumble movement, characterized by randomly reoriented ballistic movements. Understanding the mean chord length can provide foundational insights into the geometry and scale of motion within a domain, which is indirectly relevant to estimating escape times or paths in narrow escape scenarios, especially when considering the impact of domain geometry on escape dynamics Paoluzzi, Angelani, and Puglisi (2020); Rupprecht, Bénichou, and et al. (2015); Singer, Schuss, and et al. (2006) or target searching Rupprecht, Bénichou, and Voituriez (2016).
@@
It is widely recognized that numerous stochastic transport phenomena observed in the real world do not exhibit jump lengths that are exponentially distributed, nor do they show scattering angles that are uniformly random Davis and Marshak (2004); Barthelemy, Bertolotti, and Wiersma (2008); Svensson et al. (2014). For instance, isotropic random flights having fixed length have applications in polymer chains Flory (1969), and an exponential law has applications, as mentioned, to neutron diffusion processes and locomotion problems in biology. It naturally raises the question of whether these processes still adhere to similar universal characteristics. In fact, heuristic reasoning indicates that Cauchy’s formula is applicable to any (isotropic) random walk as long as the random walk’s entry into the domain involves a length distribution that aligns with equilibrium Benichou et al. (2005); Blanco and Fournier (2006); Mazzolo, de Mulatier, and Zoia (2014). In particular, some attention was recently given in the literature to the role of boundary conditions in influencing the absorption time of run-and-tumble particles Angelani (2015); Artuso and Zamora (2024), and to the particle’s speed distribution or the spatial dimensionality Mori et al. (2020).
@@
Complex dynamics in bounded domains are occasionally linked to non-homogeneous distributions, as in the case of active particles Caprini and Marconi (2019); Volpe, Gigan, and Volpe (2014); Yang, Manning, and Marchetti (2014) or when interactions with walls are present Souzy et al. (2022). Therefore, it is natural to wonder if Cauchy universality is upheld in these situations. This is the main topic we will cover, giving examples of mechanisms causing nonuniform stationary probability distributions and breaking Cauchy universality; we will take into account both the scenario where particles move in a stochastic way inside the body (random walk) and the scenario where they follow rectilinear trajectories between consecutive collisions with the boundary (billiard). Through theoretical analysis and numerical simulations, this study examines how different boundary and reorientation conditions affect the motion’s universal properties. This research offers potential implications for understanding motion in confined systems across various scientific domains, extending beyond the classical applications of billiards and random walks to include biological and physical phenomena where such conditions are prevalent.
@@
The paper is structured as follows. In the next section, the methodology for simulating particle movement within a 2D circular space is described, including variations in step length and angular distributions under different boundary conditions. Section III presents the analysis confirming the MPLT under uniformly random angular distributions and highlights deviations with non-uniform distributions. The section IV examines path length distributions and their variance, illustrating the transition from diffusive to quasi-ballistic movements. Section V explores the radial probability distributions, showing uniformity under elastic boundary conditions and deviations with random angle conditions. In section VI, we introduce memory effects in particle reorientation and their impact on path length and radial distributions. Finally, we summarize the key findings and their implications in section VII.
II Simulations
@@
Figure 1: Trajectory of the particle inside a circle with elastic boundary. $l/R$: (a) $l/R=0.01$, (b) $l/R=0.1$, (c) $l/R=1$, (d) $l/R=10$. Figure (a) is diffusive, on the other extreme, (d) is quasi-ballistic.
@@
To replicate a simple geometric setting common to experimental realizations Rashid, Long, and Singh (2019); Long et al. (2017); Frangipane et al. (2019); Souzy et al. (2022), our code simulates a particle moving within a unit 2D circular space. In our model, the particle travels at a constant unit speed, $v=1$. Therefore, the time between successive collisions is equal to the total trajectory length, allowing us to discuss average residence time and mean path length interchangeably. In two dimensions, the mean path length, as defined in eq. (1), is expressed as follows:
@@
where $l=<\ell>$ is the mean step length and $\sigma$ is the standard deviation, (b) uniformly random, (c) fixed (as in polymer chains Flory (1969)), and (d) exponential distributed (typical of bacteria mobility Frangipane et al. (2019) and neutron scattering de Kruijf and Kloosterman (2003)),
@@
In the case of the Gaussian distribution, no cutoff was used. This means that there is a possibility that the particles move backwards if $\ell$ takes a negative value. We will analyze the statistics and the validation of the MPLT under various angular distributions: (a) a fixed angle, which can be considered as a representation of strong chemotaxis (preferred direction); (b) a Gaussian distribution, representing weak chemotaxis, since there is a preferred direction with some deviation, a similar approach is used in the work of Romanczuk and Salbreux Romanczuk and Salbreux (2015); (c) the particle is allowed to move in any direction uniformly distributed in $[0,2\pi]$, which is the typical case considered in most studies.
@@
In our simulation with elastic boundaries, when a particle takes a step that would end outside the circular boundary, we detect this by calculating the distance between the new position and the center of the circle. If this distance exceeds the radius of the circle, we consider the particle to have collided with the boundary. Upon detecting such a collision, we compute the point of intersection between the particle’s trajectory and the circle boundary. The particle’s trajectory is then reflected at this boundary point. The reflection is implemented by calculating the new direction of the particle after the collision, following the principle that the angle of incidence equals the angle of reflection. The particle is then moved in this new direction, ensuring it stays within the boundary. The distance traveled by the particle after the collision is calculated as the difference between the original step length and the distance already traveled between the particle’s previous position and the boundary. If this new step also leads the particle outside the boundary, the process is repeated until the particle’s position lies within the circle after the reflection. In the case of random outgoing angle, after determining the collision point, the particle is assigned a new random direction after the collision. This is done by drawing a new angle uniformly from the range $[0,2\pi]$.
@@
By adjusting the mean step length, $l$, we can observe different types of movement, as illustrated in Fig. (1). While these examples start from the center of the circle, the findings presented in this work are not dependent on the starting location, including points on the boundary or any other random initial positions within the circle. During simulations with extremely few steps, different initial distributions could potentially influence the results, although we verified that the long-term behavior, particularly in the case of many collisions, is less sensitive to the initial distribution of walkers. In Fig. (1), the exit angle after each tumble event is uniformly random, and the step length follows an exponential distribution. We observe various movement types according to the ratio $l/R$, where $R$ is the circle’s radius. When $l/R<<1$, the movement is diffusive, but when $l/R>>1$, the movement becomes quasi-ballistic. Purely ballistic movement occurs in the limit as $l/R\to\infty$. In our simulations, a combination of both types of movements is always present. We refer to these limit movements as diffusive and quasi-ballistic based on the qualitative behavior of the system: for small $l/R$ values, the predominant movement resembles a random walk, whereas when $l/R$ is large, the predominant movement resembles a billiard (where the particle bounces inside the circle following straight paths).
III Mean path length theorem
@@
Figure 2: Difference between mean path length according to the theorem and the simulation versus the number of collisions, n, with elastic boundary condition, exponential step length distribution, and uniformly random angular distribution. Each point is an average of 20 simulations. The black solid line is a power-law fitting of the points.
First, we explore the conditions under which the Mean Path Length Theorem (MPLT) is valid. To do this, we calculate the mean path length using various combinations of step length and angular reorientation distribution. This analysis is conducted with elastic boundary conditions for four distinct values of $l/R$: $0.01$, $0.1$, $1$, and $10$. We find that the MPLT is valid only when the angular distribution is uniformly random, across all step length distributions. At first, the fact that the mean path length is the same for different values of $l/R$ seems very surprising, because if we replace straight segments (or arcs) with irregular paths, we intuitively expect a twofold mechanism to profoundly change the dynamics (stochastic short returns to the boundary and long walks in the domain without touching the boundary) Blanco and Fournier (2006). This fact is confirmed by plotting the ratio $(<L>_{s}/<L>_{t})-1$, where $<L>_{s}$ represents the mean path length from the simulation, and $<L>_{t}$ is the theoretical value predicted by the MPLT. If the MPLT is accurate, the simulated and theoretical mean path lengths will converge as the number of collisions increases. Here, we present some of the most illustrative cases.
@@
In Fig. (2), a typical graph of $(<L>_{s}/<L>_{t})-1$ is shown for various $l/R$ ratios, using an exponential step length distribution and a uniformly random angular distribution. It is observed that the difference between simulated and theoretical values tends to zero as the number of collisions increases. The step length distribution can be adjusted to any of the mentioned scenarios, and the MPLT will still hold true. For illustration, in Figs. (3(a)) and (3(b)), we present two distinct cases of step length distribution: (a) random between 00 and $2l$, and (b) a constant value of $l$. The differences between the simulated and theoretical values of mean path length approach zero as the number of collisions tends towards infinity in all instances.
@@
Figure 3: Difference between theoretical mean path length and simulation versus the number of collisions with elastic conditions at the boundary, random angle distribution, and (a) random step length between 0 and $2l$, (b) constant step length $l$. Black lines: power-law fittings.
As illustrated in the figures 2, 3(a) and 3(b), the quantity $(<L>_{s}/<L>_{t})-1)$ decreases following a power-law as the number of collisions increases. The black line in the figures are power-law fittings. According to the Central Limit Theorem, under appropriate conditions, the distribution of the sample mean converges to a normal distribution Montgomery and Runger (2014). This holds even if the original variables themselves are not normally distributed. We expect that the speed of convergence is at least of the order $1/\sqrt{N}$: indeed in our case this is not a rigorous statement, since we did not study in detail eventual correlations between successive steps.
@@
Figure 4: Difference between theoretical mean path length simulation versus the number of collisions with elastic conditions at the boundary, exponential step length distribution, and (a) gaussian angle distribution with mean 0, and $\sigma=\pi/2$; (b) fixed value of angle at $\pi/4$.
On the other hand, if we alter the distribution of the reorientation angle after the tumble events, the assumption of homogeneous flux at the boundary no longer applies, and consequently, the mean path length does not converge to the same value for all $l/R$ ratios, as illustrated in Figs. (4(a)) and (4(b)). In Fig. (4(a)), the angles after the tumble event follow a Gaussian distribution with a mean of 00 and a standard deviation of $\pi/2$. We observed similar outcomes with various combinations of means and standard deviations. Similarly, Fig. (4(b)) displays results for a fixed angle of $\pi/4$, yet comparable results are achieved with different preferred direction values. The figures suggest that the difference between the simulated and theoretical mean path length is greater when the parameter $l/R$ is small, i.e., in the diffusive regime. A diffusive regime is characterized by many small, local changes in position. Initially, the drift component acts to "straighten out" the overall trajectory, but when the boundary is reached, it causes the particle to remain in a specific region of the domain, as seen in Fig. 5, introducing significant anisotropy. In contrast, in a ballistic regime, particles move in nearly straight lines over longer distances between collisions, with little deviation in their trajectory. When a drift component is introduced into a ballistic system, the directional bias in the movement is less influential, because, from time to time, the particle takes a step long enough to escape from the preferred region of the domain, reducing the anisotropy. This leads to a smaller deviation from the MPLT predictions. The MPLT assumes a certain level of isotropy in the way particles sample the space, and as a result, the average path length measured from simulations deviates more significantly in the diffusive+drift case than in the ballistic+drift case.
@@
Figure 5: Trajectory of 1000 steps of the particle inside a circle with elastic boundary, exponential step length distribution, and $l/R=0.1$. (a) The angle distribution normal-distributed around $\theta=0$, (b) Angle distribution fixed at $\pi/4$.
The fixed angle and the Gaussian distribution around an angle create a preferred direction, observable in some experimental behaviors of bacteria due to mechanisms like chemotaxis. The fixed angle constitutes a strong condition (the particle consistently moves in a specific direction), while the Gaussian distribution represents a less stringent condition (there is a preferred direction, yet a certain probability exists for the particle to move in another direction close to the preferred one). In scenarios with fixed and Gaussian-distributed angles, the particle tends to collide with the boundary consistently in the same region, as illustrated in Fig. (5). This results in non-uniformity and anisotropy, thus violating the assumptions underlying the MPLT. This outcome is not surprising, as previous works have already clarified the conditions of the angular distribution under which the theorem holds Bardsley and Dubi (1981); Mazzolo (2004); Benichou et al. (2005); Mazzolo, de Mulatier, and Zoia (2014).